OpenStudy (anonymous):

How would I rewrite the following expression as a single logarithm? (Writing expression in comments)

OpenStudy (anonymous):

2log(x+3) - 3log(x-7) + 5log(x-2) - log(x^2)

OpenStudy (anonymous):

I was given options to choose from -- I'm not asking for a straightup answer if nobody wants to give me one, just help on how to do it. A.\[\log(\frac{ (x+3)^{2}(x-2)^{5} }{ (x-7)^{3}x ^{2} })\]B.\[\log(\frac{ x ^{2}(x+3)^{2}(x-7)^{3} }{ (x-2)^{5} })\]C.\[\log(\frac{ (x+3)^{2}(x-7)^{3} }{ x ^{2}(x-2)^{5} })\]D.\[\log(\frac{ (x+3)^{2}(x-7)^{3} }{ (x-2)^{-5}x ^{-2} }\]

OpenStudy (anonymous):

First have you narrowed it down to two answers? deductive reasoning?

OpenStudy (anonymous):

Now that I really think about it the only answer that makes sense to me is A

OpenStudy (anonymous):

good u solved on ur own YAY

OpenStudy (anonymous):

haha yay